A chef is using a mixture of two brands of Italian dressing. The first brand contains 10% vinegar, and the second brand contains 15% vinegar. The chef wants to make 260 milliliters of a dressing that is 13% vinegar. How much of each brand should she use?

Answer by mananth(16949)

A chef is using a mixture of two brands of Italian dressing. The first brand contains 10% vinegar, and the second brand contains 15% vinegar. The chef wants to make 260 milliliters of a dressing that is 13% vinegar. How much of each brand should she use?

let 10% vinegar required be x
15% vinegar required will be 260-x
0.1x+0.15(260-x)= 260*13
0.1x+39-0.15x= 33.8
-005x=-39 +33.8
x= 5.2/-0.05
x=104 milliters which is 10% brand
260-104=56 which is 15% brand